Session: Nov 1998 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina ( at Fri Nov 27 22:32:17 MET 1998) Minor failures

Medicina (A. Cattani at Fri Nov 27 22:33:05 MET 1998) Success

Onsala (John Conway at Sat Nov 28 18:00:57 MET 1998) Minor failures
No known problems

Torun (K.Borkowski at Mon Nov 30 10:23:51 MET 1998) Success
A few seconds earlier tape stops at the low tape sensor. No other problems reported.

Noto (C.Trigilio at Mon Nov 30 11:16:20 MET 1998) Success

Effelsberg (dgraham at Mon Nov 30 15:10:52 MET 1998) Minor failures
seems ok

Jodrell (Alastair Gunn at Tue Dec 1 08:11:19 MET 1998) Minor failures
Elevation servo failure at 19:07, restarted at 19:17. Field System Log file for this experiment is missing - investigating.

Westerbork (Tony Foley at Tue Dec 1 18:27:24 MET 1998) Minor failures
Problems with the very short scans on 1954+513, with the array missing some of them until we fiddled with the integration time parameters

Calibration (Jean\-Francois Desmurs at Fri Dec 11 16:36:22 MET 1998) Minor failures
Missing values for JB (missing log file) The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: gb031.antab. Read the comments in the file to get further information.